OpenStudy (anonymous):

The equation 5x^2-10x+5=0 can be written in the form (px+q)^2+r=0, where p,q and r are integers. Find the value of p,q and r.

OpenStudy (anonymous):

\[5(x-1)^2=0\]

OpenStudy (anonymous):

how, did you equate like terms?

OpenStudy (anonymous):

actually no. i factored out the 5 form the expression to get \[5(x^2-2x+1)=0\] and then i recognize \[x^2-2x+1\] as a perfect square, namely \[x^2-2x+1=(x-1)^2\] so we get \[5(x-1)^2=0\]

OpenStudy (anonymous):

what are the values of p,q and r

OpenStudy (anonymous):

p^2x^2=5x^2 so p^2=5 so p=??????

OpenStudy (anonymous):

actually i don't think we can write this as \[(px+q)^2+r=0\] but we can write this one \[(x-1)^2=0\] in that form. \[p=1, q =-1, r = 0\]

OpenStudy (anonymous):

really?